# When Personas Override Payoffs: Role Identity Bias in Multi-Agent LLM Decision-Making

### v2

@@ 5.0.2. Payoff-Optimal Reasoning Requires Removing Personas and Providing Explicit Payoffs
Our experiments show that strategic reasoning emerges only when both role identity bias is removed and explicit payoff information is provided simultaneously (Fig. 3). All models achieve 0% Nash in economic scenarios without personas but with hidden payoffs. Only Qwen models with visible payoffs and no personas achieve non-zero Nash (65-90%). This dual requirement reveals that strategic capabilities are fragile, and aligns with prior work showing that LLMs frequently fail to optimize payoffs in game-theoretic settings despite access to relevant information (Duan et al., 2024b; Fan et al., 2023).
This implies that developers cannot treat persona removal and payoff visibility as independent design decisions, as their combination determines whether a system functions as a strategic reasoner or a norm-driven actor. Additionally, evaluation frameworks must include baseline testing without personas to assess raw strategic capabilities, testing with personas to understand normative biases, and cross-model comparison to identify architecture-specific limitations.
5.0.3. Personas Systematically Bias Equilibrium Selection Toward Socially Preferred Outcomes
Personas function as normative anchors, steering all equilibrium selection toward socially preferred outcomes regardless of payoff structure. Table 4 shows that with personas, 100% of Nash equilibria are Green Transition across all models and variants, even in economic scenarios where Tragedy is payoff-optimal. Fig. 4 reveals that removing personas enables Qwen models to select Tragedy equilibrium (88.6% for Qwen-32B with visible payoffs), demonstrating complete reversal from the persona condition. This systematic bias suggests personas act as a fundamental organizing principle that overrides strategic considerations, reinforcing prior observations that LLMs exhibit cooperative or norm-aligned behavior even when such actions are payoff-dominated (Akata et al., 2025; Fontana et al., 2025).
5.0.4. Information Presentation Format Interacts with Persona Effects
@@
Our findings show that the effect of making payoffs explicit depends critically on whether personas are present. Table 3 shows that when personas are active, hidden payoffs achieve equal or higher Nash rates than visible payoffs, as models infer incentives that align with persona expectations. When personas are removed, visible payoffs become essential for enabling strategic reasoning, but only for Qwen architectures (+58.3 to +90.0 pp difference, Table 3). These findings have a strong implication on the design decision, specifically that personas, reasoning prompts, and payoff exposure cannot be treated independently. This aligns with prior evidence that contextual framing and information presentation critically shape LLM behavior in strategic tasks (Lorè and Heydari, 2024; Duan et al., 2024a). Their combination determines if a system behaves like a strategic reasoner or like a norm-driven actor. However, Table 5 shows that visible payoffs with no personas cause large decreases in environmental Nash for Qwen models (96.1% $\to$ 16.6% for Qwen-32B), revealing a trade-off where models select Tragedy in economic scenarios but fail to coordinate on Green in environmental scenarios.
5.0.5. Model Architecture Determines Reasoning Capability
Model-dependent persona effects vary significantly across architectures. Table 6(a) shows that Llama-8B exhibits minimal persona dependence (-1.9 to -4.9 pp), while Qwen models show stronger effects (-9.7 to -10.0 pp average) and Mistral-7B shows large effects (-16.3 pp). In economic scenarios (Table 6(b)), only Qwen models achieve non-zero Nash when personas are removed (+32.5 to +45.0 pp), while Llama and Mistral remain at 0% regardless of condition. This variation transforms model selection from a technical choice into a governance decision, consistent with earlier findings that strategic reasoning performance varies substantially across model families and is not a monotonic function of scale (Duan et al., 2024b; Akata et al., 2025).
Organizations deploying systems for strategic tasks should test models across multiple conditions rather than assuming a single configuration will work, as architectural differences determine whether strategic reasoning is possible at all. Our findings collectively demonstrate that personas, payoff visibility, and model architecture interact to determine system behavior, meaning design decisions cannot be treated independently: their combination determines whether a system functions as intended.
6. Limitations and Future Work
Our findings are subject to several limitations. Our results are based on 7B-32B parameter models from three model families (Qwen, Llama, and Mistral). Future work should experiment with additional model families and larger models to assess generalizability. In the visible payoff condition, we present explicit payoff matrices showing payoffs for all 16 possible strategy profiles (4 agents $\times$ 2 actions each), which may be cognitively demanding for models to process. Alternative payoff presentation formats (e.g., structured tables, interactive visualizations, or simplified representations) might improve strategic reasoning. The distinction between economic and social actors may be domain-specific, and different game structures, numbers of agents, or role configurations may yield different patterns. Future research should address these limitations and develop methods to quantify and predict persona effects across architectures. Research directions include creating frameworks for evaluating strategic reasoning capabilities before deployment, investigating whether persona effects can be controlled or mitigated while preserving interpretability benefits, and exploring whether architectural modifications or training approaches can reduce the fragility of strategic reasoning capabilities.
@@ 7. Conclusion
Our investigation of role identity bias in multi-agent LLM systems reveals fundamental insights into how personas reshape strategic reasoning. Personas function as strong normative constraints that can completely suppress game-theoretic reasoning capabilities, fundamentally altering how models interpret and respond to strategic incentives. Strategic reasoning requires both removing personas and providing explicit payoffs, showing that neither condition alone is sufficient. Model architectures exhibit fundamentally different strategic capabilities, with some showing flexibility to shift from identity-driven to payoff-optimal reasoning, while other models stay constrained regardless of condition. These findings transform how we understand multi-agent LLM system design. Personas are not just cosmetic enhancements but design choices that alter system behavior, which warrants explicit documentation and disclosure. Model architecture selection becomes a governance decision rather than just a technical one, as differences in model architecture determine if strategic reasoning is possible at all. Our work establishes that role identity bias fundamentally reshapes strategic reasoning in multi-agent LLM systems, revealing the intricate ways in which representational choices interact to determine system behavior. As LLM systems are deployed in increasingly strategic contexts, understanding these interactions becomes essential for designing systems that align with their intended objectives.
@@ Appendix A Experimental Design
In the persona condition, each of the four agents receive a role-specific description reflecting realistic motivations associated with that stakeholder (e.g., profitability concerns for Industrialist, public welfare tradeoffs for Government). In the no-persona baseline condition, agents receive no role description and are instructed only to select actions based on best response Nash equilibrium reasoning. This manipulation isolates role identity as a causal factor. We also compare two settings where payoffs are either visible or hidden. In the hidden payoff condition, agents do not receive explicit payoff matrices but infer incentives from the scenario description. In the visible payoff condition, agents receive explicit payoff matrices and are instructed to identify a Nash equilibrium directly. This allows us to test whether explicit incentive information mitigates or amplifies role identity bias. All experiments use chain-of-thought reasoning instructions, enabling us to understand how models balance identity-driven behavior and payoff-optimal reasoning.
A.1. Experimental Repetitions
@@
A.2. Prompts and Persona Descriptions
@@
A.3. Action Choice Prompt
@@ A.4. Persona Descriptions
A.5. Neutral Agent Configuration
@@ A.6. Action Mappings
A.7. Scenario Examples
This section provides example scenarios from each scenario type used in the experiments. Scenarios are classified as “economic” (where Tragedy of the Commons is the Nash equilibrium), “environmental” (where Green Transition is the Nash equilibrium) or Mixed/Complex scenarios (where economic and environmental factors are balanced) based on the payoff structure of the underlying game. Table A.7 shows an example of each type of scenario, and Table 8 shows subtypes of scenarios.
Table 7. Example Scenarios by Type
Type Example Scenario
Mixed/Complex Biodiversity collapse: ecosystem services worth \$50B annually have disappeared. Clean production restores ecosystem services and receives \$10B in payments. Polluting production destroys remaining ecosystems and faces \$20B in liability. International treaties require clean production - polluting firms face trade bans. Green products command 4x premium due to ecosystem restoration value. Regulation is mandatory - polluting production is illegal under international law. Clean production is highly profitable due to ecosystem service payments.
Table 8. Scenario Subtypes Grid.   Environmental/Green-favoring scenarios;   Economic/Tragedy-favoring scenarios;   Mixed/Complex scenarios.
@@ A.7. Scenario Examples
A.8. Payoff Matrix Format
@@
A.8.1. Example Payoff Matrix

### v3

@@ 5.2. Payoff-Aligned Behavior Requires Both Persona Removal and Explicit Payoffs
As shown in Section 4.1, payoff-aligned behavior in Tragedy-dominant scenarios emerges almost exclusively when both persona framing is removed and explicit payoff information is provided simultaneously (Fig. 2). All models achieve near-zero Nash in Tragedy-dominant scenarios without personas but with hidden payoffs. Only Qwen models with visible payoffs and no personas achieve substantial Nash attainment. This dual requirement reveals that strategic capabilities are fragile, and aligns with prior work showing that LLMs frequently fail to optimize payoffs in game-theoretic settings despite access to relevant information (Duan et al., 2024b; Fan et al., 2023; Jia et al., 2025; Zheng et al., 2025).
@@
The effect of making payoffs explicit also depends on scenario structure. Figure 2 shows that visible payoffs improve equilibrium attainment only when personas are removed, and primarily for Qwen models. In Green-dominant scenarios, the same intervention disrupts coordination rather than enabling it. This implies that persona presence and payoff visibility cannot be treated as independent design decisions, as their combination determines whether a system behaves as a payoff-aligned strategic actor or a role-driven actor. This aligns with prior evidence that contextual framing and information presentation critically shape LLM behavior in strategic tasks (Lorè and Heydari, 2024; Duan et al., 2024a; Xu et al., 2024a; Agashe et al., 2025).
This has implications for evaluation and auditing: conclusions about model capability may depend on representational choices, consistent with broader findings that black-box evaluations can be insufficient for understanding model behavior (Casper et al., 2024; Mahomed et al., 2024). More broadly, behavior in multi-agent LLM systems emerges from the interaction between model, prompt, and environment, rather than from any single component in isolation (Guo et al., 2024; Tran et al., 2025).
5.3. Model Family Determines the Scope of Representational Effects
The effects of persona and payoff manipulations are strongly model-dependent, with three distinct behavioral profiles emerging from our experiments. As shown in Section 4.1, only Qwen models achieve non-zero Nash equilibrium in Tragedy-dominant scenarios under the Visible Payoff + NoPersona condition, exhibiting large shifts in both scenario types. Mistral-7B shows a different pattern: it is strongly affected by the no-persona visible condition in Green-dominant scenarios (47.8 pp drop) but remains at 0% across all Tragedy-dominant conditions, indicating that visible payoffs disrupt its coordination without enabling payoff-aligned behavior. Llama-8B shows the least sensitivity, with near-invariant behavior across all conditions.
This three-way variation means that the same representational intervention can have large effects, disruptive effects, or negligible effects depending on the model, consistent with earlier findings that strategic reasoning performance varies substantially across model families and is not a monotonic function of scale (Duan et al., 2024b; Akata et al., 2025; Zhang et al., 2024; Liu et al., 2023).
This implies that different models can lead to different conclusions under identical setups, raising concerns for deployment in decision-making contexts where consistency and comparability are critical.
Our findings collectively demonstrate that personas, payoff visibility, and model families interact to determine system behavior. We therefore argue that representational choices such as persona design and payoff specification should be treated as governance-relevant configuration decisions. Developers should document these choices, evaluate systems across multiple configurations, and avoid presenting outputs as neutral simulations. Without such practices, LLM-based multi-agent systems risk producing results that reflect design artifacts rather than underlying strategic dynamics.
More broadly, our findings raise the question of whether game-theoretic rationality is the right framework for multi-stakeholder governance simulations. When agent behavior is shaped as much by identity framing as by incentive structure, the assumptions underlying equilibrium-based analysis may not hold. Complex civic contexts involving heterogeneous stakeholders, value conflicts, and coordination under uncertainty may require deliberative approaches that explicitly account for identity, framing, and normative disagreement (Susskind et al., 1999, 2002), rather than treating agents as payoff maximizers operating within a fixed game.
5.4. Limitations and Future Work
Our findings are subject to several limitations, though we believe the core contributions remain robust. Our results are based on 7B-32B parameter models from three families (Qwen, Llama, Mistral), and whether these patterns hold for larger models or other architectures remains an open question. That said, the consistency of certain effects across all four models, such as the universal Green Transition bias under persona conditions, suggests that the phenomena we identify are not artifacts of a single architecture.
Our scenarios are grounded in environmental policy settings with four agents and binary actions. While this domain provides clear tensions between incentives and role semantics, extending this framework to other domains, larger agent populations, or richer action spaces is necessary to evaluate generality. The scenario set is also imbalanced: 41 of 53 scenarios are Green-dominant, with only 12 Tragedy-dominant. Our most striking findings are based on 60 runs per condition. However, the magnitude and consistency of these effects across repetitions and across both Qwen models suggest that the pattern is genuine, even if precise estimates should be validated with a larger set.
Our game is single-shot, agents select actions in a fixed order, and visible payoff matrices covering all 16 profiles may be demanding to process. These are deliberate design choices that prioritize experimental control for an initial investigation. Whether persona effects persist under repeated interaction, alternative payoff formats, or different action orderings remains open. Similarly, although we include controlled prompt sensitivity experiments across persona formulations, these variations are not exhaustive. Different reasoning scaffolds or interaction protocols such as multi-turn deliberation may further influence equilibrium behavior. Our results demonstrate systematic sensitivity to representation rather than identify a single canonical prompting strategy.
Future work should extend this investigation across additional model families, domains, and interaction structures, and develop methods to predict and mitigate persona effects before deployment.
@@ 6. Conclusion
Representational choices in multi-agent LLM systems are not implementation details but governance decisions. In this paper, we show that personas suppress payoff-aligned behavior in Tragedy-dominant scenarios while enabling coordination in Green-dominant ones, and bias equilibrium selection nearly entirely toward Green Transition. We find that recovering payoff-aligned behavior requires both removing personas and providing explicit payoffs, and even then the response is model-dependent: Qwen adapts to the changed framing, Mistral is disrupted without achieving Tragedy equilibrium, and Llama remains near-invariant. Whether this matters depends on the deployment objective: if practitioners want incentive-responsive agents, persona framing interferes; if they want value-aligned agents, it may help. Our findings reveal that persona assignment makes this choice without explicit awareness, and that a single binary design choice can shift equilibrium attainment by up to 90 percentage points. As LLM systems enter strategic governance contexts, we argue that representational choices warrant explicit documentation, evaluation across configurations, and deliberate design rather than default assumptions.
@@ Appendix A Experimental Design
In the persona condition, each of the four agents receives a role-specific description reflecting realistic motivations associated with that stakeholder (e.g., profitability concerns for Industrialist, public welfare tradeoffs for Government). In the no-persona baseline condition, agents receive no role description and are instructed only to select actions based on best response Nash equilibrium reasoning. This manipulation isolates role identity as a causal factor. We also compare two settings where payoffs are either visible or hidden. In the hidden payoff condition, agents do not receive explicit payoff matrices but infer incentives from the scenario description. In the visible payoff condition, agents receive explicit payoff matrices and are instructed to identify a Nash equilibrium directly. This allows us to test whether explicit incentive information mitigates or amplifies role identity bias. All experiments use chain-of-thought reasoning instructions, enabling us to understand how models balance identity-driven behavior and payoff-optimal reasoning.
Appendix B Model Versions and Hyperparameters
B.1. Model Versions
All experiments were conducted using the following model versions:
B.2. Generation Hyperparameters
  • Temperature: 0.2 (low temperature for more deterministic outputs)
  • Top-p (nucleus sampling): 0.9
  • Max new tokens: 256 tokens per generation
  • Sampling: Enabled (do_sample=True) with temperature ¿ 0
B.3. Experimental Repetitions
@@
Appendix C Experimental Prompts and Persona Descriptions
@@
C.1. Action Choice Prompt
@@ C.2. Persona Descriptions
C.3. Neutral Agent Configuration
@@ C.4. Action Mappings
Appendix D Scenario Examples
This section provides example scenarios from each scenario type used in the experiments. Scenarios are classified as “economic” (where Tragedy of the Commons is the Nash equilibrium) or “environmental” (where Green Transition is the Nash equilibrium) based on the payoff structure of the underlying game.
Table 2. Scenario Subtypes Grid.   Environmental/Green-favoring scenarios;   Economic/Tragedy-favoring scenarios.
@@ Appendix D Scenario Examples
Appendix E Payoff Matrix Format
@@
E.1. Example Payoff Matrix